# Privacy Vulnerabilities in Marginals-based Synthetic Data

### v1

@@ A-D Complexity and runtime
TABLE VII: Runtimes of MAMA-MIA on SNAKE data when $\varepsilon=10$ (i.e. experiment B), separated into the shadow modelling stage, where we select FPs, and the inference stage, where we use them to compute $\zeta$. The runtime of the select FP stage is for 50 shadow modelling runs. The runtime of the compute $\zeta$ stage given is the average runtime for one trial, attacking all targets in $D_{target}$. Results for Private-GSD on 31,623 records are omitted because of computational limitations.
MAMA-MIA on MST MAMA-MIA on PrivBayes MAMA-MIA on Private-GSD
100 13.2 min 2.57 sec 2.11 sec 6.73 sec 2.57 min 16.3 sec
316 13.3 min 3.85 sec 2.63 sec 5.33 sec 2.69 min 38.8 sec
1,000 14.0 min 4.14 sec 6.46 sec 4.74 sec 4.93 min 79.9 sec
3,162 14.3 min 3.75 sec 15.7 sec 5.22 sec 8.09 min 2.17 min
10,000 14.5 min 3.02 sec 50.0 sec 4.37 sec 10.2 min 2.52 min
31,623 41.7 min 1.01 sec 4.37 min 4.43 sec N/a N/a
@@
The second stage is where we perform the density estimation to calculate membership predictions for $D_{target}$. For each focal-point in MST and Private-GSD, we construct a marginal probability table for $D_{aux}$, $D_{synth}$. We make the reasonable assumption of the ability to use amortized constant-time hash tables during this construction, and so constructing these takes linear time with respect to the size of the datasets, at most size $n$.
@@
Then for each record in $D_{target}$, we look up the probability of its value. Together, these steps for one focal-point amount to $\mathcal{O}(|D_{aux}|+|D_{synth}|+2|D_{target}|)$ time, which is just $\mathcal{O}(n)$, since $D_{target}\subseteq D_{aux}$, and assuming $|D_{synth}|<|D_{aux}|$. This is also the case for PrivBayes, where we build conditional probability tables instead. Like marginal tables, constructing conditional tables with hash tables requires looking at each feature in the focal-point for each record once, since the number of features in each conditional is practically small. So the total runtime of calculating $\zeta$ is $\mathcal{O}(|\mathcal{F}|\cdot n)$ where $|\mathcal{F}|$ is the amount of focal-points chosen by the SDG algorithm during shadow modelling.
Runtimes are given in Table VII for this “compute $\zeta$” stage separately. PrivBayes takes slightly longer than MST, which is likely due to the fact that it chooses conditionals as focal-points, which offer higher specificity than MST’s marginals proper, and so presumably stress the hashing capability. Computing $\zeta$ in Private-GSD appears to take much longer than on the other two SDG algorithms. This can be explained by $|\mathcal{F}|$ being much larger in Private-GSD, which selects thousands of focal-points over one-hot encoded data, as opposed to the dozen or so FPs on un-encoded data chosen by MST and PrivBayes. The reason why the time to compute $\zeta$ is relatively constant, despite the growing dataset size, is because the $\zeta$ also performs computations over the entire $D_{aux}$, which is much larger (i.e. over 200,000 rows) than any $D_{synth}$ tried, and so dominates that portion of the runtime.
@@
A-E Extended results
Main results in greater detail We offer AUC results in Tables V, and VI for experiment A, conducted on the SNAKE data and the housing data. These tables include MAMA-MIA and TAPAS accuracies, averaged over 30 trials with standard deviations. These correspond to Figures 3 and 5 respectively. The runtimes of MAMA-MIA for this experiment, now separated into its shadow modelling and its density estimation step, are given in Table VII.
Showing the impact of shadow modelling in MAMA-MIA Shadow modelling to select focal-points (FPs) is a time consuming step in MAMA-MIA, especially when attacking synthetic data generated by SDGs that are in themselves computationally expensive to train (see Table VII). One may wonder what the effect is of this careful selection of focal-points on the accuracy of the attack, i.e. whether it is worth the computational effort.
Figure 15: MAMA-MIA and DOMIAS-using-KDE accuracies when performing set MI on SNAKE data, where $|D_{target}|=1,000$ records, $|D_{target}|=32$ sets, and $|D_{member}|=16$ sets.
@@ A-E Extended results
Figure 16: MAMA-MIA and DOMIAS-using-KDE accuracies when performing set MI on housing data, where $|D_{target}|=1,000$ records, $|D_{target}|=32$ sets, and $|D_{member}|=16$ sets.
@@
Set MI experiments We show results for set MI experiments in Figures 15 and 16. These are conducted using the same settings as in experiment A, i.e. with $|D_{train}|$ fixed at 1,000 records, trialling both the SNAKE and the housing data. We only present results for DOMIAS, when using KDE, and we omit set MI results for the attack from the TAPAS library because it does not offer set MI in its API. Sets are groups of at least five target records, and instead of making $|D_{target}|$, and $|D_{member}|$ equal to 32 and 16 records, respectively, we use 32 and 16 sets. The results are unsurprising, where MAMA-MIA’s accuracy improves over the single MI setting.
A-F Limitations

### v2

@@ A-D Complexity and runtime
The second stage is where we perform the density estimation to calculate membership predictions for $D_{target}$. For each focal-point in MST and Private-GSD, we construct a marginal probability table for $D_{aux}$, $D_{synth}$. We make the reasonable assumption of the ability to use amortized constant-time hash tables during this construction, and so constructing these takes linear time with respect to the size of the datasets, at most size $n$. Then for each record in $D_{target}$, we look up the probability of its value. Together, these steps for one focal-point amount to $\mathcal{O}(|D_{aux}|+|D_{synth}|+2|D_{target}|)$ time, which is just $\mathcal{O}(n)$, since $D_{target}\subseteq D_{aux}$, and assuming $|D_{synth}|<|D_{aux}|$. This is also the case for PrivBayes, where we build conditional probability tables instead. Like marginal tables, constructing conditional tables with hash tables requires looking at each feature in the focal-point for each record once, since the number of features in each conditional is practically small. So the total runtime of calculating $\zeta$ is $\mathcal{O}(|\mathcal{F}|\cdot n)$ where $|\mathcal{F}|$ is the amount of focal-points chosen by the SDG algorithm during shadow modelling.
Runtimes are given in Table V for this “compute $\zeta$” stage separately. PrivBayes takes slightly longer than MST, which is likely due to the fact that it chooses conditionals as focal-points, which offer higher specificity than MST’s marginals proper, and so presumably stress the hashing capability. Computing $\zeta$ in Private-GSD appears to take much longer than on the other two SDG algorithms. This can be explained by $|\mathcal{F}|$ being much larger in Private-GSD, which selects thousands of focal-points over one-hot encoded data, as opposed to the dozen or so FPs on un-encoded data chosen by MST and PrivBayes. The reason why the time to compute $\zeta$ is relatively constant, despite the growing dataset size, is because the $\zeta$ also performs computations over the entire $D_{aux}$, which is much larger (i.e. over 200,000 rows) than any $D_{synth}$ tried, and so dominates that portion of the runtime.
Figure 12: ROC curves of MIAs in experiment A when $\varepsilon=10$.
A-E Extended results
Main results in greater detail We offer AUC results in Tables VI, and VII for experiment A, conducted on the SNAKE data and the housing data. These tables include MAMA-MIA and TAPAS accuracies, averaged over 30 trials with standard deviations. These correspond to Figures 3 and 5 respectively. The runtimes of MAMA-MIA for this experiment, now separated into its shadow modelling and its density estimation step, are given in Table V.
Additionally, we include full ROC curves in Figure 12 of the TAPAS attack, DOMIAS, and MAMA-MIA for experiment A, when $\varepsilon=10$. To average across all trials, we simply compute the curve using all 960 target predictions (across the 30 trials, with 32 targets per trial). These results show in more detail the accuracies of MAMA-MIA predictions at different false positive rates.
Showing the impact of shadow modelling in MAMA-MIA Shadow modelling to select focal-points (FPs) is a time consuming step in MAMA-MIA, especially when attacking synthetic data generated by SDGs that are in themselves computationally expensive to train (see Table V). One may wonder what the effect is of this careful selection of focal-points on the accuracy of the attack, i.e. whether it is worth the computational effort.
@@ A-E Extended results
Figure 16: MAMA-MIA and DOMIAS-using-KDE accuracies when performing set MI on the SNAKE data (top) and on the housing data (bottom). $|D_{target}|=1,000$ records, $|D_{target}|=32$ sets, and $|D_{member}|=16$ sets.
@@
Set MI experiments We show results for set MI experiments in Figure 16. These are conducted using the same settings as in experiment A, i.e. with $|D_{train}|$ fixed at 1,000 records, trialling both the SNAKE and the housing data. We only present results for DOMIAS, when using KDE, and we omit set MI results for the attack from the TAPAS library because it does not offer set MI in its API. Sets are groups of at least five target records, and instead of making $|D_{target}|$, and $|D_{member}|$ equal to 32 and 16 records, respectively, we use 32 and 16 sets. The results are unsurprising, where MAMA-MIA’s accuracy improves over the single MI setting.
MAMA-MIA’s performance on the wrong synthetic data generator Table VIII shows the interesting case when MAMA-MIA is tailored to one SDG algorithm, say MST, and is used perform membership inference on data produced by a different SDG algorithm. We show all combinations of tailoring MAMA-MIA to one of MST, PrivBayes, and Private-GSD, and using it to attack data from each of those SDGs. The columns depict which algorithm generated the data, and the rows depict which algorithm MAMA-MIA was tailored to. Naturally, the diagonal (in boldface) represents the case where MAMA-MIA attacks the correct data. This scenerio is carried out on the SNAKE data, when $\varepsilon=10$ and $|D_{train}|=1,000$.
These results are interesting because they show that MAMA-MIA performs better than random guessing when the data is not from the anticipated generator. This is despite using focal-points from an entirely different SDG algorithm in our construction of $\zeta$. However, as expected, MAMA-MIA attacking the data for which it is tailored produces far better results, demonstrating the advantage MAMA-MIA has with knowledge of which algorithm was used.
TABLE VIII: AUC scores $\pm$ one standard deviation, using MAMA-MIA tailored to one SDG algorithm to perform inference on synthetic data produced by another algorithm. $\varepsilon=10$, using SNAKE data, and $|D_{train}|=1,000$. The diagonal values (in boldface) are when MAMA-MIA attacks data generated by the intended SDG algorithm.
Attack method Data generated by
MST PrivBayes Private-GSD
MAMA-MIA $\mathbf{0.614\pm.11}$ $0.504\pm.13$ $0.523\pm.10$
MST
MAMA-MIA $0.539\pm.11$ $\mathbf{0.578\pm.09}$ $0.545\pm.08$
PrivBayes
MAMA-MIA $0.546\pm.11$ $0.538\pm.10$ $\mathbf{0.612\pm.10}$
Private-GSD
A-F Description of related MIAs
Groundhog attack[34]: Stadler et al. assess the privacy gain of different SDGs by proposing targeted attacks. They consider a scenario in which only one copy of the generated synthetic data is published. They describe their threat model as a generic black-box MIA. Specifically, they consider an adversary who has prior knowledge of the SDG (referred to as black-box knowledge in our paper) and access to a reference dataset (referred to as auxiliary data in our paper) that shares the same distribution as the training data. The underlying intuition of their attack is that if the “real training data” contains certain attributes, the synthetic data is likely to preserve these attributes, such as correlations and statistical metrics, even if the SDG model does not explicitly capture them.
To perform this attack, the adversary employs a shadow modeling strategy following Shokri et al. [33]. The core idea is to sample a large number of training datasets from the auxiliary dataset, which are then subsequently and independently used to generate synthetic datasets. The goal of the adversary is then to infer if a specific target $t$ was included in the training set used to generate the synthetic dataset being attacked. To achieve this, Stadler et al. propose to train a Random Forest classifier with a rich set of input features – (a) naive features such as median, mean and variance of each column in tabular dataset, (b) histogram based features that represent distributions of each column in tabular dataset, and (c) correlation features that capture the correlation coefficients. For further details, please refer to [34] and its implementation in TAPAS [^21] https://tapas-privacy.readthedocs.io/en/latest/library-of-attacks.html
TAPAS [17]: TAPAS is a comprehensive framework comprising various types of attacks to evaluate the privacy of generated synthetic data across different scenarios, including different threat models and privacy evaluation metrics. In their paper, Houssiau et al. make novel tweaks to the GroundHog attack above to improve the accuracy of these attacks. Specifically, they adjust the set of features provided to the Random Forest Classifier. They randomly select $x$ subsets of $m$-ordered column feature sets in the tabular data. For each subset, they then compute the number of records in the synthetic data that match the target record $t$ on the selected column feature sets. These $x$ queries form the dimensions of the resulting feature vector that is fed to the Random Forest Classifier  [^22] We implemented this improved attack using the RandomTargetedQueryFeature in TAPAS library and extended it to consider a set of target records rather than a single target.
DOMIAS [39]: DOMIAS proposed a MIA on synthetic data that leverages density estimation to detect if the generator was overfit on the training dataset. The authors assume a modified black-box setting where the adversary has access to one copy of synthetic dataset, a target record $t$ and additional access to a reference dataset (auxiliary dataset in our paper). DOMIAS constructs the density estimators on the reference ${p}_{ref}()$ and the synthetic ${p}_{syn}()$ datasets independently and then follows the equation below for scoring the MIA for target $t$
$\mathcal{A_{DOMIAS}}=\frac{{p_{syn}(t)}}{{p_{ref}(t)}}$
The intuition is similar to what we explain in section III and Figure 2.
A-G Limitations